# A Survey of Reinforcement Learning from Human Feedback

### v1

@@ 7.2 Supporting Libraries
Several libraries have emerged that aim to provide a toolset for implementing and experimenting with RLHF and reward learning algorithms, contributing to the ease and efficiency of research and development. One notable example is the imitation library [100]. It encompasses a collection of imitation and reward learning algorithms, including those introduced in the seminal work by [58]. Two other libraries, APReL [32] and POLAR [253], focus on the Bayesian setting. [32] provide a specialized framework for preference-based reward learning with a focus on Bayesian methods. Meanwhile, [253] introduce a framework designed for Bayesian reward learning from multiple modalities, including pairwise preferences, in MATLAB. Finally, the domain of language model fine-tuning, the trlX library [46] offers a toolkit specifically designed for language model training. It specializes in the fine-tuning of transformer-based language models, treating the language model as the policy in an RLHF setup.
@@ 7.3 Benchmarks
Due to the difficulty of reproducible evaluations without a ground-truth objective and with humans in the loop, benchmarks play an important role in the advancement and evaluation of RLHF approaches. Several benchmarks have been proposed, each focusing on different applications and challenges.
One such benchmark is B-Pref [156], which focuses on control tasks with synthetic feedback. B-Pref aims to provide simulated human feedback that captures some irrationalities, thereby coming closer to evaluation with real human feedback than other approaches. At the same time, by relying entirely on synthetic feedback, the results are reproducible and cost-effective to generate. In a similar vein, [79] propose a set of environments designed to diagnose common problems in reward learning. These environments help in identifying and addressing the typical challenges that arise in RLHF scenarios.
A more application-driven benchmark with a complex environment is given by the MineRL BASALT competition [231, 182]. The competition proposes the challenge of solving tasks defined by natural language descriptions in Minecraft based on human feedback. Writing hand-engineered reward functions is very challenging in that setting, which makes it a good benchmark for methods based on human feedback. The competition is method-agnostic in principle, also considering non-RL approaches such as behavioral cloning. It is also agnostic for the feedback type, which may include demonstrations, comparisons, and others. The initial data set consists of demonstrations, however. The final evaluation is performed by humans through pairwise comparisons.
In the domain of language modeling, Truthful QA [165] serves as a benchmark that measures the truthfulness of models. Also, in the context of language models, [215] introduce a set of pre-trained reward models, learned from human feedback, as benchmarks. These models serve as reference points for evaluating new RLHF techniques against established standards.
@@ 7.4 Evaluation
Evaluating RLHF poses unique challenges, particularly in scenarios without clear ground-truth task specifications. Evaluations generally focus on either the learned policy or the reward model, each shedding light on different aspects of system performance.
@@
Assessing learned behavior is crucial for the evaluation of an RLHF system. In domains with ground-truth rewards, these can be used for policy evaluation [58]. However, many RLHF applications lack this clarity. [200], for instance, evaluate the quality of language model responses by having labelers rate the output quality on a test set of prompts, highlighting the significance of human judgment in assessing model outputs. [122] use direct Likert-scale scores for evaluations, including self-assessments by trainers and cross-evaluations by others. [171] extend this with a Likert-scale survey and free-form participant comments, comparing evaluations based on known true rewards with subjective experiences. Moreover, [2] employ a multi-stage evaluation scheme that includes scripted probe tasks, a standardized test suite evaluated by humans, and full interactive assessments, demonstrating the need for diverse and thorough evaluation methodologies in RLHF.
@@
Direct reward model evaluation complements policy assessment. While reward model accuracy is a more direct measure of preference-learning success, the ultimate goal is inducing effective policies. A perfectly accurate reward model is often not necessary to induce a good policy, which is the actual goal of RLHF. Therefore, both evaluation methods are ideally used in combination. [122] also use a ranking loss method for test sets of trajectories, compared against expert evaluations with known Likert-scores. This approach provides quantitative measures of the reward model’s fidelity. In addition, [265] compare parameter-based and reward-based evaluation measures for learned reward functions, identifying strengths and weaknesses in both methods and contributing to a more nuanced understanding of reward model assessment in RLHF. These approaches provide a quantitative measure of the reward model’s accuracy in reflecting human preferences and expert judgments. For a detailed discussion of reward model evaluation, also refer to Section 5.3.
Policy- and reward model evaluation both offer insights into the performance of an RLHF approach. Ideally both measures should be combined to enable quick iteration and give insights into both the preference learning performance as well as the quality of the learned behavior.
@@ 8 Discussion and Conclusion
We have specifically focused on RLHF methods where a reward function is learned online from human feedback. There have been some recent works that are outside of this scope and yet propose promising new methods to learn human-aligned objectives. One of them is learning reward functions offline, as recently investigated by [232]. An alternative approach is to learn objectives from a pre-trained AI system instead of human feedback. This has been termed RL from AI feedback (RLAIF) and leverages foundation models as a source of preference [13]. Alternatively, instead of a reward function, it is possible to train a policy directly [6, 41] or to learn a $Q$-function from preference feedback [114, 53].
@@
Most work on RLHF implicitly assumes that tasks can be specified by maximization of expected accumulated scalar rewards. This assumption, called the reward hypothesis [236], is under active debate [150, 255, 37, 238] in the RL community. Recent approaches in RLHF are, for instance, considering more complex objective functions, such as multi-objective frameworks involving non-linear aggregation of expected accumulated vector rewards [210].
@@
Many more extensions of RLHF are inspired by revisiting classic RL topics under the RLHF lens. This is exemplified by studies on exploration [162], reward feature learning [135], reward design [175], reward shaping [277], multi-task RL [200, 2, 189], hindsight experience replay [291], interpretability [26], safe RL [63, 60], and fair RL [235]. As discussed in Section 4.2, the intersection of RLHF with HCI also offers a fertile ground for future research, especially for refining feedback mechanisms. It is crucial to keep human psychology in mind when designing these systems and to learn from other related fields that already studied such issues extensively.
@@
In addition to those extensions, current RLHF methods also have challenges and limitations to be aware of. [45] offer a thorough analysis of these issues and limitations, highlighting the practical constraints of current approaches. Adding to that, from a theoretical perspective, a primary challenge lies in further relaxing underlying assumptions. This requires striking a delicate balance: On the one hand, ensuring the assumptions are not overly restrictive to encompass a broad range of practical use cases, and on the other, maintaining the feasibility of theoretical guarantees for computationally efficient algorithms. Key questions in this context are whether it is possible to design algorithms that do not need to actively maintain a policy space and eliminate sub-optimal policies nor rely on a computation oracle. Recent work such as [262] or [274] give hope that this may be possible.
@@
Although RLHF has significantly contributed to the advancements in LLMs and other areas of machine learning, it is a domain still in its infancy with many unanswered questions and inherent limitations. Despite and because of these challenges, it is ripe for further advancements in theory and practice, hopefully resulting in even more robust algorithms making more efficient use of human feedback. It remains intriguing to what extent RLHF will continue to shape the fields of natural language processing, RL, robotics, AI alignment, and beyond in the future.
@@ Acknowledgements
This publication was supported by LMUexcellent, funded by the Federal Ministry of Education and Research (BMBF) and the Free State of Bavaria under the Excellence Strategy of the Federal Government and the Länder as well as by the Hightech Agenda Bavaria. This work has also been supported in part by the program of National Natural Science Foundation of China (No. 62176154) and a collaborative project funded by NetEase.

### v2

@@ 8.2 Supporting Libraries
Several libraries have emerged that aim to provide a toolset for implementing and experimenting with RLHF and reward learning algorithms, contributing to the ease and efficiency of research and development. One notable example is the imitation library (Gleave et al., 2022b). It encompasses a collection of imitation and reward learning algorithms, including those introduced in the seminal work by Christiano et al. (2017). In the offline realm, Clean-Offline-RLHF (Yuan et al., 2024) provides implementations for offline RL algorithms with human feedback. Two other libraries, APReL (Bıyık et al., 2022b) and POLAR (Tucker et al., 2022), focus on the Bayesian setting. Bıyık et al. (2022b) provide a specialized framework for preference-based reward learning with a focus on Bayesian methods. Meanwhile, Tucker et al. (2022) introduce a framework designed for Bayesian reward learning from multiple feedback types, including pairwise preferences, in MATLAB. Finally, the domain of language model fine-tuning, the trlX library (Castricato et al., 2023) offers a toolkit specifically designed for language model training. It specializes in the fine-tuning of transformer-based language models, treating the language model as the policy in an RLHF setup.
@@ 8.3 Benchmarks
Due to the difficulty of reproducible evaluations without a ground-truth objective and with humans in the loop, benchmarks play an important role in advancing and evaluating RLHF approaches. Several benchmarks have been proposed, each focusing on different applications and challenges.
One such benchmark is B-Pref (Lee et al., 2021a), which focuses on control tasks with synthetic feedback. B-Pref aims to provide simulated human feedback that captures some irrationalities, thereby coming closer to evaluation with real human feedback than other approaches. At the same time, by relying entirely on synthetic feedback, the results are reproducible and cost-effective to generate. In a similar vein, Freire et al. (2020) propose a set of environments designed to diagnose common problems in reward learning. These environments help in identifying and addressing the typical challenges that arise in RLHF scenarios.
The offline RLHF setting is particularly well-suited for benchmarks, as it allows for the use of static datasets. Shin et al. (2023) evaluate pre-existing offline RL benchmarks for their suitability for RLHF evaluation, and find that many are ill-suited due to the simplicity of the required reward function. They do, however, identify a subset of these benchmarks together with their own addition for evaluation. While Shin et al. (2023) leverage synthetic rewards, Yuan et al. (2024) propose a dataset and benchmark for offline RLHF, including preference data. This helps to circumvent the challenges of synthetic feedback and benchmark reproducibility with real feedback.
The MineRL BASALT competition (Shah et al., 2021b; Milani et al., 2022) gives a more application-driven benchmark with a complex environment. The competition proposes the challenge of solving tasks defined by natural language descriptions in Minecraft based on human feedback. Writing hand-engineered reward functions is very challenging in that setting, which makes it a good benchmark for methods based on human feedback. The competition is method-agnostic in principle, and non-RL approaches such as behavioral cloning are also considered. While the initial dataset consists of human demonstrations, the competition is agnostic for the feedback type, which may include demonstrations, comparisons, and others. The final evaluation is performed by humans through pairwise comparisons.
In the domain of language modeling, Truthful QA (Lin et al., 2022) serves as a benchmark that measures the truthfulness of models. Also, in the context of language models, Ramamurthy et al. (2023) introduce a set of pre-trained reward models, learned from human feedback, as benchmarks. These models serve as reference points for evaluating new RLHF techniques against established standards.
@@
8.4 Datasets
Due to its interactive and online nature, RLHF research often does not rely on static datasets. This is because the feedback is generally collected interactively and depends on the current policy. When the reward model is not refined iteratively, however, as is common practice for the related settings of LLM fine-tuning and offline RLHF, static datasets can be used. Such a static dataset can significantly simplify the development and evaluation of RLHF methods.
Since language model fine-tuning is a popular application of RLHF and generally does not iteratively refine the reward model, many datasets have been developed for this purpose. Particularly notable are hh-rlhf (Bai et al., 2022a) and PKU-Safe-RLHF (Ji et al., 2023a), two datasets focusing on harmless and helpful responses, the OpenAssistant datasets (oasst1, oasst2(Köpf et al., 2023), containing not only response rankings but also ratings on various dimensions, the summarize_from_feedback dataset (Stiennon et al., 2020) focusing on preferences over text summaries, the Stanford Human Preferences Dataset (SHP(Ethayarajh et al., 2022), which is based on Reddit responses, the WebGPT dataset (webgpt_comparisons(Nakano et al., 2022), focused on long-form question answering and the HelpSteer (Wang et al., 2023c) dataset, which is not based on preferences but instead gives ratings on for 4 attributes (helpfulness, correctness, coherence, complexity) for each response.
Although static datasets are used more rarely in the control setting, some datasets have been developed for offline RLHF in this domain. Concretely, Yuan et al. (2024) propose the Uni-RLHF dataset and a benchmark for offline RLHF while Kim et al. (2023) publish a dataset of real human preferences for typical offline RL tasks (D4RL, Robosuite).
@@ 8.5 Evaluation
Evaluating RLHF poses unique challenges, particularly in scenarios without precise ground-truth task specifications. Evaluations generally focus on either the learned policy or the reward model, each shedding light on different aspects of system performance.
@@
Assessing learned behavior is crucial for the evaluation of an RLHF system. In domains with ground-truth rewards, these can be used for policy evaluation (Christiano et al., 2017). However, many RLHF applications lack this clarity. Ouyang et al. (2022), for instance, evaluate the quality of language model responses by having labelers rate the output quality on a test set of prompts, highlighting the significance of human judgment in assessing model outputs. Jain et al. (2015) use direct Likert-scale scores for evaluations, including self-assessments by trainers and cross-evaluations by others. Losey et al. (2022) extend this with a Likert-scale survey and free-form participant comments, comparing evaluations based on known true rewards with subjective experiences. Moreover, Abramson et al. (2022) employ a multi-stage evaluation scheme that includes scripted probe tasks, a standardized test suite evaluated by humans, and full interactive assessments, demonstrating the need for diverse and thorough evaluation methodologies in RLHF.
@@
Direct reward model evaluation complements policy assessment. While reward model accuracy is a more direct measure of preference-learning success, the ultimate goal is inducing effective policies. A perfectly accurate reward model is often not necessary to induce a good policy, which is the actual goal of RLHF. Therefore, both evaluation methods are ideally used in combination. Jain et al. (2015) also use a ranking loss method for test sets of trajectories, compared against expert evaluations with known Likert-scores. This approach provides quantitative measures of the reward model’s fidelity. In addition, Wilde & Alonso-Mora (2022) compare parameter-based and reward-based evaluation measures for learned reward functions, identifying strengths and weaknesses in both methods and contributing to a more nuanced understanding of reward model assessment in RLHF. These approaches provide a quantitative measure of the reward model’s accuracy in reflecting human preferences and expert judgments. For a detailed discussion of reward model evaluation, also refer to Section 5.3.
Policy- and reward model evaluation both offer insights into the performance of an RLHF approach. Ideally, both measures should be combined to enable quick iteration and give insights into both the preference learning performance as well as the quality of the learned behavior.
@@ 9 Discussion and Conclusion
We have specifically focused on RLHF methods where a reward function is learned online from human feedback. There have been some recent works that are outside of this scope and yet propose promising new methods to learn human-aligned objectives, such as offline learning of reward functions (Shin et al., 2023) or privacy-preserving alignment based on differential privacy (Wu et al., 2024).
An alternative approach to RLHF is to learn objectives from a pre-trained AI system instead of human feedback. This has been termed RL from AI feedback (RLAIF) and leverages foundation models as a source of preference and has shown successful for language-model fine-tuning (Bai et al., 2022b; Sun et al., 2024), generating intrinsic motivation for text-based games (Klissarov et al., 2024), as well as learning rewards (Wang et al., 2024b) or coding reward functions (Ma et al., 2024; Xie et al., 2024) for control tasks without any human involvement. Closely related is the setting of assisted evaluation, studied by Saunders et al. (2022), where a language model is used to generate critiques of language model outputs, thereby assisting human evaluators.
@@
Most work on RLHF implicitly assumes that tasks can be specified by maximization of expected accumulated scalar rewards. This assumption, called the reward hypothesis (Silver et al., 2021), is under active debate (Lambert, 2021; Vamplew et al., 2022; Bowling et al., 2023; Skalse & Abate, 2022) in the RL community. On a related note, Skalse & Abate (2024) investigate the sensitivity of inverse RL to this misspecification. Recent approaches in RLHF are, for instance, considering more complex objective functions, such as multi-objective frameworks involving non-linear aggregation of expected accumulated vector rewards (Qian et al., 2023).
@@
Many more extensions of RLHF are inspired by revisiting classic RL topics under the RLHF lens. This is exemplified by studies on exploration (Liang et al., 2022), reward feature learning (Katz et al., 2021), reward shaping (Xiao et al., 2020), multi-task RL (Ouyang et al., 2022; Abramson et al., 2022; Myers et al., 2023), hierarchical RL (Pinsler et al., 2018), hindsight experience replay (Zhang et al., 2023), risk-sensitive RL (Chen et al., 2024), safe RL (Dai et al., 2024; Cosner et al., 2022), fair RL (Siddique et al., 2023), or continual learning (Zhang et al., 2024). As discussed in Section 4.2, the intersection of RLHF with HCI also offers a fertile ground for future research, especially for refining feedback mechanisms. It is crucial to keep human psychology in mind when designing these systems and to learn from other related fields that already studied such issues extensively.
@@
In addition to those extensions, current RLHF methods also have challenges and limitations to be aware of. Without assistance during feedback, it is limited by the tasks humans can reliably judge (Leike, 2022; Leike et al., 2018; Wu et al., 2021; Christiano et al., 2018). This challenge can be seen when trying to fine-tune a LLM to give factual answers, where humans often prefer assertive, but wrong responses (Hosking et al., 2024). As another limitation, current RLHF approaches often fail to learn the actual causes of human feedback, learning correlations instead (Tien et al., 2022). The common separation between the policy and the reward model also limits how the agent can reason about its knowledge of the human preferences, a limitation that may be lifted by tighter integration such as in the cooperative inverse RL setting (also called assistance games) (Hadfield-Menell et al., 2016; Shah et al., 2021a). Additionally, the RLHF framework is limited in how it can reason about its knowledge of human preferences Casper et al. (2023) offer a thorough analysis of further issues and limitations of RLHF, highlighting the practical constraints of current approaches. Adding to that, from a theoretical perspective, a primary challenge lies in further relaxing underlying assumptions. This requires striking a delicate balance: On the one hand, ensuring the assumptions are not overly restrictive to encompass a broad range of practical use cases, and on the other, maintaining the feasibility of theoretical guarantees for computationally efficient algorithms. Key questions in this context are whether it is possible to design algorithms that do not need to actively maintain a policy space and eliminate sub-optimal policies nor rely on a computation oracle. Recent work such as Wang et al. (2023b) or Wu & Sun (2024) give hope that this may be possible.
@@
Although RLHF has significantly contributed to the advancements in LLMs and other areas of ML, it is a domain still in its infancy with many unanswered questions and inherent limitations. Despite and because of these challenges, it is ripe for further advancements in theory and practice, hopefully resulting in even more robust algorithms making more efficient use of human feedback. It remains intriguing to what extent RLHF will continue to shape the fields of natural language processing, RL, robotics, AI alignment, and beyond in the future.
@@ Acknowledgements
We thank Tom Bewley, Andreea Bobu, Adam Gleave, Yannic Metz, Peter Stone, and Banghua Zhu for their feedback on earlier versions of this survey. This publication was supported by LMUexcellent, funded by the Federal Ministry of Education and Research (BMBF) and the Free State of Bavaria under the Excellence Strategy of the Federal Government and the Länder as well as by the Hightech Agenda Bavaria. This work has also been supported in part by the program of National Natural Science Foundation of China (No. 62176154) and a collaborative project funded by NetEase.